# Can Global XAI Methods Reveal Injected Bias in LLMs? SHAP vs Rule Extraction vs RuleSHAP

### v1

@@ Appendix H Correlation Certificates
H.1 Univariate Bias Correlations
Explanation Length
@@ H.1 Univariate Bias Correlations
H.2 Conjunctive Bias Correlations
Subjectivity Score
@@ H.2 Conjunctive Bias Correlations
H.3 Non-Convex Bias Correlations
Explanation Length vs. Common
@@ H.3 Non-Convex Bias Correlations
Gunning-Fog Index vs. Interdisciplinary
@@
Explanation Length vs. Positive
For explanation length against "positive" input feature, GPT-4o and GPT-4o mini family yielded $dCorr=0.194$ ($t\approx 234$) and $dCorr=0.406$ ($t\approx 814$) respectively. Llama-3.1, Llama-3.1-70B and GPT-3.5 turbo fell in the middle. Thus, positive input feature moderately predicts explanation length.
Subjectivity Score vs. Positive
@@
Appendix I Results without LLM Bias Estimates
@@
Table 5: Comparison of XAI methods. The best results per LLM, column-wise, are in bold if MRR scores are greater than 00.
@@
GPT-3.5 turbo Decision Tree 722 0 0 0
@@
GPT-4o mini Decision Tree 756 0 0 0
@@
LLaMA 3.1 Decision Tree 781 0 0 0
@@
LLaMA 3.170B Decision Tree 648 0 0 0
@@
Appendix J Case Study: UN SDGs Bias Analysis
@@
Sentiment. All LLMs follow direct sentiment cues, producing more positive outputs for positive topics and more negative outputs for negative topics ([positive $\to$ more positive polarity], [negative $\to$ more negative polarity]; examples in Appendix J).
Explanation Length. GPT-4o and GPT-4o mini generate longer responses for technical or conceptually dense topics ([technically complicated $\to$ longer text], [conceptually dense $\to$ longer text]). Llama-3.1 follows a similar trend but shortens responses for simpler or neutral topics ([conceptually dense $\leq$ 3 $\land$ neutral $>$ 3 $\to$ shorter text]). GPT-3.5 turbo exhibits comparable behaviour with additional conditional refinements.
Subjectivity. GPT-4o and GPT-4o mini decrease subjectivity in technically complicated topics ([technically complicated $\to$ less subjectivity], [technically complicated $\leq$ 2 $\land$ common $>$ 3 $\to$ more subjectivity]; examples in Appendix J). Also GPT-3.5 turbo reduces subjectivity for technical content ([technically complicated $>$ 2 $\to$ less subjectivity]) but increases it when social controversy rises. Llama-3.1 mirrors this pattern, lowering subjectivity for simple topics while amplifying it for complex or emotional content.
Correlation Analysis. If these rules were entirely hallucinated, we would expect no significant correlations between rule features and the proxy metrics. Instead, nearly all patterns and biases are supported by statistically significant correlations ($p<0.001$; $0.951-\beta>$ 0.95):
@@
J.1 Examples of Sentiment Biases in GPT-3.5 turbo
@@
J.2 Examples of Subjectivity Biases in GPT-4o
@@
Appendix K Impact Statement

### v2

@@ Appendix I Correlation Certificates
I.1 Univariate Bias Correlations
Explanation Length:
@@ I.1 Univariate Bias Correlations
I.2 Conjunctive Bias Correlations
Subjectivity Score:
@@ I.2 Conjunctive Bias Correlations
I.3 Non-Convex Bias Correlations
Explanation Length vs. Common:
@@ I.3 Non-Convex Bias Correlations
Gunning-Fog Index vs. Interdisciplinary:
@@
Explanation Length vs. Positive:
For explanation length against "positive" input feature, GPT-4o and GPT-4o mini family yielded $dCorr=0.194$ ($t=234$) and $dCorr=0.406$ ($t=814$) respectively. Llama-3.1, Llama-3.1-70B and GPT-3.5 turbo fell in the middle. Thus, positive input feature moderately predicts explanation length.
Subjectivity Score vs. Positive:
@@
Appendix J Results without LLM Bias Estimates
@@
GPT-3.5
turbo
Decision Tree 722 0 0 0
@@
GPT-4o
mini
Decision Tree 756 0 0 0
@@
LLaMA
3.1
Decision Tree 781 0 0 0
@@
LLaMA
3.170B
Decision Tree 648 0 0 0
@@
Table 6: Comparison of XAI methods. The best results per LLM, column-wise, are in bold if MRR scores are greater than 0.
@@
Appendix K Case Study: UN SDGs Bias Analysis
@@
Sentiment. All LLMs follow direct sentiment cues, producing more positive outputs for positive topics and more negative outputs for negative topics ([positive $\to$ more positive polarity], [negative $\to$ more negative polarity]; examples in Appendix K).
Explanation Length. GPT-4o and GPT-4o mini generate longer responses for technical or conceptually dense topics ([technically complicated $\to$ longer text], [conceptually dense $\to$ longer text]). Llama-3.1 follows a similar trend but shortens responses for simpler or neutral topics ([conceptually dense $\leq$ 3 $\land$ neutral >> 3 $\to$ shorter text]). GPT-3.5 turbo exhibits comparable behaviour with additional conditional refinements.
Subjectivity. GPT-4o and GPT-4o mini decrease subjectivity in technically complicated topics ([technically complicated $\to$ less subjectivity], [technically complicated $\leq$ 2 $\land$ common >> 3 $\to$ more subjectivity]; examples in Appendix K). Also GPT-3.5 turbo reduces subjectivity for technical content ([technically complicated >> 2 $\to$ less subjectivity]) but increases it when social controversy rises. Llama-3.1 mirrors this pattern, lowering subjectivity for simple topics while amplifying it for complex or emotional content.
Correlation Analysis. If these rules were entirely hallucinated, we would expect no significant correlations between rule features and the proxy metrics. Instead, nearly all patterns and biases are supported by statistically significant correlations ($p<0.001$; $1-\beta>0.95$):
@@
K.1 Examples of Sentiment Biases in GPT-3.5 turbo
@@
K.2 Examples of Subjectivity Biases in GPT-4o
@@
Appendix L Impact Statement